a shopkeeper sold 4 times as many eggs on monday than on tuesday. if the total number of eggs on these two days was 850. find the number of eggs sold(a) monday and (b) on tuesday

Answer by Cromlix(4381)

Hi there,
Ratio 1:4
Add together 1 + 4 = 5
This is your denominator
Monday = 1/5 x 850
= 170 eggs
Tuesday = 4/5 x 850
= 680 eggs
Hope this helps:-)